# Heights and periods of algebraic cycles in families

### v1

@@ 4. Proof of the finiteness result à la Ullmo
Before moving on to the proof, let us take a step back and look at this proposition from a historical point of view. In studying the André–Oort conjecture, Ullmo [Ull14, Thm. 4.1] proved this finiteness result for maximal weakly special subvarieties (a particular kind of weakly optimal subvarieties) in the case of pure Shimura varieties as an application of the pure Ax–Lindemann theorem (a special case of Ax–Schanuel). The finiteness is ultimately obtained by the following fact: any countable set which is definable in an o-minimal structure is finite.
@@
Recently, Baldi–Urbanik [BU24, Thm. 7.1] proved Proposition 3.8 for general admissible VMHS in a different way (but also uses Ax–Schanuel as a core). Weakly optimal subvarieties are called monodromically atypical maximal, and Proposition 3.8 is called Geometric Zilber–Pink as in [BKU24]. Their proof does not use o-minimality, and gives some effective results.
@@ 4.1. Zariski optimal subsets and an application of mixed Ax–Schanuel
A complex analytic irreducible subset $\tilde{Y}$ of $u^{-1}(S)$ is called Zariski optimal if the following holds true: $\tilde{Y}\subsetneq\tilde{Y}^{\prime}\subseteq u^{-1}(S)\Rightarrow\delta_{\mathrm{Zar}}(\tilde{Y})<\delta_{\mathrm{Zar}}(\tilde{Y}^{\prime})$, with $\tilde{Y}^{\prime}$ complex analytic irreducible.
@@
Assume $\tilde{Y}\subseteq u^{-1}(S)$ is Zariski optimal. Then $\tilde{Y}^{\mathrm{Zar}}=\tilde{Y}^{\mathrm{ws}}$ and $\tilde{Y}$ is a complex analytic irreducible component of $\tilde{Y}^{\mathrm{ws}}\cap u^{-1}(S)$.
@@
Let $\tilde{Y}\subseteq u^{-1}(S)$ be Zariski optimal. Then $\tilde{Y}$ is a complex analytic irreducible component of $\tilde{Y}^{\mathrm{Zar}}\cap u^{-1}(S)$. Hence it suffices to prove $\tilde{Y}^{\mathrm{Zar}}=\tilde{Y}^{\mathrm{ws}}$.
Let $\tilde{Y}^{\prime}$ be a complex analytic irreducible component of $\tilde{Y}^{\mathrm{ws}}\cap u^{-1}(S)$ which contains $\tilde{Y}$. Then $u(\tilde{Y}^{\prime})$ is closed algebraic in $S$ because $u(\tilde{Y}^{\mathrm{ws}})\cap S$ is a closed algebraic subset of $S$ by Proposition 3.3.
Assume $\tilde{Y}\not=\tilde{Y}^{\prime}$. Then $\delta_{\mathrm{Zar}}(\tilde{Y})<\delta_{\mathrm{Zar}}(\tilde{Y}^{\prime})$ by the Zariski optimality of $\tilde{Y}$. So
$\dim\tilde{Y}^{\mathrm{Zar}}-\dim\tilde{Y}<\dim\tilde{Y}^{\prime\mathrm{Zar}}-\dim\tilde{Y}^{\prime}\leq\dim\tilde{Y}^{\mathrm{ws}}-\dim\tilde{Y}^{\prime}.$
Now that $u(\tilde{Y})\subseteq u(\tilde{Y}^{\prime})$, we have $u(\tilde{Y})^{\mathrm{Zar}}\subseteq u(\tilde{Y}^{\prime})$ because the right hand side is closed algebraic in $S$. So
$\dim u(\tilde{Y})^{\mathrm{Zar}}\leq\dim u(\tilde{Y}^{\prime})=\dim\tilde{Y}^{\prime}.$
@@
$\dim\tilde{Y}^{\mathrm{Zar}}+\dim u(\tilde{Y})^{\mathrm{Zar}}<\dim\tilde{Y}^{\mathrm{ws}}+\dim\tilde{Y}.$
This contradicts the weak mixed Ax–Schanuel theorem (Theorem 3.5) for $\tilde{Y}$.
Hence we must have $\tilde{Y}=\tilde{Y}^{\prime}$. In particular, $u(\tilde{Y})$ is algebraic. So $\tilde{Y}^{\mathrm{Zar}}=\tilde{Y}^{\mathrm{ws}}$. ∎
For $Y\subseteq S$ irreducible, let $\tilde{Y}$ be a complex analytic irreducible component of $u^{-1}(Y)$.
@@
If $Y\subseteq X$ is weakly optimal, then $\tilde{Y}$ is Zariski optimal in $u^{-1}(S)$.
@@
Let $\tilde{Y}^{\prime}\supseteq\tilde{Y}$ be a complex analytic irreducible subset of $u^{-1}(S)$ such that $\delta_{\mathrm{Zar}}(\tilde{Y})\geq\delta_{\mathrm{Zar}}(\tilde{Y}^{\prime})$. We may and do assume that $\tilde{Y}^{\prime}$ is Zariski optimal.
Set $Y^{\prime\prime}:=u(\tilde{Y}^{\prime})^{\mathrm{Zar}}$. Then $Y^{\prime\prime}\subseteq u(\tilde{Y}^{\prime,\mathrm{ws}})\cap S$ because $u(\tilde{Y}^{\prime,\mathrm{ws}})\cap S$ is closed and algebraic. Thus
$\delta_{\mathrm{ws}}(Y^{\prime\prime})=\dim Y^{\prime\prime,\mathrm{ws}}-\dim Y^{\prime\prime}\leq\dim\tilde{Y}^{\prime,\mathrm{ws}}-\dim\tilde{Y}^{\prime}.$
Since $\tilde{Y}^{\prime}$ is Zariski optimal, we have $\tilde{Y}^{\prime,\mathrm{Zar}}=\tilde{Y}^{\prime,\mathrm{ws}}$ by Theorem 4.2. So, the inequality above further implies
$\delta_{\mathrm{ws}}(Y^{\prime\prime})\leq\dim\tilde{Y}^{\prime,\mathrm{Zar}}-\dim\tilde{Y}^{\prime}=\delta_{\mathrm{Zar}}(\tilde{Y}^{\prime})\leq\delta_{\mathrm{Zar}}(\tilde{Y})\leq\dim\tilde{Y}^{\mathrm{ws}}-\dim\tilde{Y}=\delta_{\mathrm{ws}}(Y).$
So $Y^{\prime\prime}=Y$ because $Y$ is weakly optimal. Thus $\tilde{Y}=\tilde{Y}^{\prime}$ is Zariski optimal. ∎
@@ 4.2. A parametrization of Zariski optimal subsets
Fix a Levi decomposition $\mathbf{G}=V\rtimes\mathbf{G}_{0}$. Then for the identification ${\mathscr{D}}=V({\mathbb{R}})\times{\mathscr{D}}_{0}$ in Proposition B.3.(i), the action of $\mathbf{G}({\mathbb{R}})$ on ${\mathscr{D}}$ is given by $(v,g_{0})\cdot(x_{V},x_{0})=(v+g_{0}x_{V},g_{0}x_{0})$.
@@
    $\{\tilde{Y}^{\mathrm{Zar}}=\tilde{Y}^{\mathrm{ws}}:\tilde{Y}\subseteq u^{-1}(S)\text{ Zariski optimal, with }\tilde{Y}\cap\mathfrak{F}\not=\emptyset\}$.
@@
Take $(x,H)\in{\mathscr{N}}^{0}$. We wish to prove that $H({\mathbb{R}})x$ equals $\tilde{Y}^{\mathrm{Zar}}$ for some Zariski optimal $\tilde{Y}\subseteq u^{-1}(S)$ which passes through $x\in\mathfrak{F}$.
Let $\tilde{Y}^{\prime}$ be a complex analytic irreducible component of $u^{-1}(S)\cap H({\mathbb{R}})x$ which passes through $x$ with $\dim\tilde{Y}^{\prime}=d(x,H)$.
Take $\tilde{Y}\supseteq\tilde{Y}^{\prime}$ with $\tilde{Y}\subseteq u^{-1}(S)$ complex analytic irreducible and $\delta_{\mathrm{Zar}}(\tilde{Y})\leq\delta_{\mathrm{Zar}}(\tilde{Y}^{\prime})$. We may and do assume that $\tilde{Y}$ is Zariski optimal. Then $\tilde{Y}^{\mathrm{Zar}}=\tilde{Y}^{\mathrm{ws}}$ is a weak Mumford–Tate domain by Theorem 4.2, and hence $\tilde{Y}^{\mathrm{Zar}}=H^{\prime}({\mathbb{R}})x$ for some $(x,H^{\prime})\in{\mathscr{N}}$. Thus $\delta_{\mathrm{Zar}}(\tilde{Y})=\delta(x,H^{\prime})$.
Thus $\tilde{Y}^{\prime}$ is contained in some irreducible component of $H({\mathbb{R}})x\cap H^{\prime}({\mathbb{R}})x$, which by Lemma 4.4.(ii) is $H^{\prime\prime}({\mathbb{R}})x$ for some $(x,H^{\prime\prime})\in{\mathscr{N}}$. Hence $d(x,H)=d(x,H^{\prime\prime})$, so by definition of ${\mathscr{N}}^{0}$ we have $H({\mathbb{R}})x=H^{\prime\prime}({\mathbb{R}})x$. So $H({\mathbb{R}})x\subseteq H^{\prime}({\mathbb{R}})x$. But $\delta(x,H)=\delta(x,H^{\prime})$. So by definition of ${\mathscr{N}}^{0}$ we have $H({\mathbb{R}})x=H^{\prime}({\mathbb{R}})x$, and hence $H({\mathbb{R}})x=\tilde{Y}^{\mathrm{Zar}}$. Now we have proved that the first set is contained in the second set and the “Moreover” part of the proposition.
Conversely let $\tilde{Y}\subseteq u^{-1}(S)$ be Zariski optimal with $\tilde{Y}\cap\mathfrak{F}\not=\emptyset$. By Theorem 4.2, $\tilde{Y}^{\mathrm{Zar}}=\tilde{Y}^{\mathrm{ws}}$ is a weak Mumford–Tate domain, and hence equals $H({\mathbb{R}})x$ for some $(x,H)\in{\mathscr{N}}$. We may furthermore choose this $x$ to be a smooth point of $H({\mathbb{R}})x\cap u^{-1}(S)\cap\mathfrak{F}$; then $\dim\tilde{Y}=d(x,H)$.
@@
Assume there exists $(x,H^{\prime})\in{\mathscr{N}}$ such that $H^{\prime}({\mathbb{R}})x\subsetneq H({\mathbb{R}})x$ and $d(x,H)=d(x,H^{\prime})$. Then $\dim\tilde{Y}=d(x,H^{\prime})$ and hence $\tilde{Y}\subseteq H^{\prime}({\mathbb{R}})x$ by analytic continuation. But then $\tilde{Y}^{\mathrm{Zar}}\subseteq H^{\prime}({\mathbb{R}})x$ by Lemma 4.4.(i). So $\tilde{Y}^{\mathrm{Zar}}\subseteq H^{\prime}({\mathbb{R}})x\subsetneq H({\mathbb{R}})x=\tilde{Y}^{\mathrm{Zar}}$, which is impossible. So for any $(x,H^{\prime})\in{\mathscr{N}}$, we have $H^{\prime}({\mathbb{R}})x\subsetneq H({\mathbb{R}})x\Rightarrow d(x,H)>d(x,H^{\prime})$.
Next, assume there exists $(x,H^{\prime})\in{\mathscr{N}}$ such that $H({\mathbb{R}})x\subsetneq H^{\prime}({\mathbb{R}})x$ and $\delta(x,H)\geq\delta(x,H^{\prime})$. Take a complex analytic irreducible component $\tilde{Y}^{\prime}$ of $u^{-1}(S)\bigcap H^{\prime}({\mathbb{R}})x$ such that $\dim\tilde{Y}^{\prime}=d(x,H^{\prime})$. Then $\tilde{Y}\subseteq\tilde{Y}^{\prime}$ by analytic continuation. By Lemma 4.4.(i), we have $\tilde{Y}^{\prime,\mathrm{Zar}}\subseteq H^{\prime}({\mathbb{R}})x$. So $\delta_{\mathrm{Zar}}(\tilde{Y}^{\prime})\leq\delta(x,H^{\prime})\leq\delta(x,H)=\delta_{\mathrm{Zar}}(\tilde{Y})$. Hence $\tilde{Y}^{\prime}=\tilde{Y}$ because $\tilde{Y}$ is Zariski optimal. So $\delta(x,H^{\prime})=\delta(x,H)$ and $d(x,H^{\prime})=d(x,H)$, and so

### v2

@@ 4. Proof of the finiteness result à la Ullmo
Before moving on to our proof, let us take a step back and look at this proposition from a historical point of view. In studying the André–Oort conjecture, Ullmo [Ull14, Thm. 4.1] proved this finiteness result for maximal weakly special subvarieties (a particular kind of weakly optimal subvarieties) in the case of pure Shimura varieties as an application of the pure Ax–Lindemann theorem (a special case of Ax–Schanuel). The finiteness is ultimately obtained by the following fact: any countable set which is definable in an o-minimal structure is finite.
@@
Recently, Baldi–Urbanik [BU24, Thm. 7.1] proved Proposition 3.8 for general admissible VMHS in a different way (but also uses Ax–Schanuel as a core). Weakly optimal subvarieties are called monodromically atypical maximal, and Proposition 3.8 is called Geometric Zilber–Pink as in [BKU24]. Their proof does not use o-minimality and gives some effective results.
@@ 4.1. Zariski optimal subsets and an application of mixed Ax–Schanuel
A complex analytic irreducible subset $\widetilde{Y}$ of $u^{-1}(S)$ is called Zariski optimal if the following holds: $\widetilde{Y}\subsetneq\widetilde{Y}^{\prime}\subseteq u^{-1}(S)\Rightarrow\delta_{\mathrm{Zar}}(\widetilde{Y})<\delta_{\mathrm{Zar}}(\widetilde{Y}^{\prime})$, with $\widetilde{Y}^{\prime}$ complex analytic irreducible.
@@
Assume $\widetilde{Y}\subseteq u^{-1}(S)$ is Zariski optimal. Then $\widetilde{Y}^{\mathrm{Zar}}=\widetilde{Y}^{\mathrm{ws}}$ and $\widetilde{Y}$ is a complex analytic irreducible component of $\widetilde{Y}^{\mathrm{ws}}\cap u^{-1}(S)$.
@@
Let $\widetilde{Y}\subseteq u^{-1}(S)$ be Zariski optimal. Then $\widetilde{Y}$ is a complex analytic irreducible component of $\widetilde{Y}^{\mathrm{Zar}}\cap u^{-1}(S)$. Hence it suffices to prove $\widetilde{Y}^{\mathrm{Zar}}=\widetilde{Y}^{\mathrm{ws}}$.
Let $\widetilde{Y}^{\prime}$ be a complex analytic irreducible component of $\widetilde{Y}^{\mathrm{ws}}\cap u^{-1}(S)$ which contains $\widetilde{Y}$. Then $u(\widetilde{Y}^{\prime})$ is closed algebraic in $S$ because $u(\widetilde{Y}^{\mathrm{ws}})\cap S$ is a closed algebraic subset of $S$ by Proposition 3.3.
Assume $\widetilde{Y}\not=\widetilde{Y}^{\prime}$. Then $\delta_{\mathrm{Zar}}(\widetilde{Y})<\delta_{\mathrm{Zar}}(\widetilde{Y}^{\prime})$ by the Zariski optimality of $\widetilde{Y}$. So
$\dim\widetilde{Y}^{\mathrm{Zar}}-\dim\widetilde{Y}<\dim\widetilde{Y}^{\prime\mathrm{Zar}}-\dim\widetilde{Y}^{\prime}\leq\dim\widetilde{Y}^{\mathrm{ws}}-\dim\widetilde{Y}^{\prime}.$
Now that $u(\widetilde{Y})\subseteq u(\widetilde{Y}^{\prime})$, we have $u(\widetilde{Y})^{\mathrm{Zar}}\subseteq u(\widetilde{Y}^{\prime})$ because the right hand side is closed algebraic in $S$. So
$\dim u(\widetilde{Y})^{\mathrm{Zar}}\leq\dim u(\widetilde{Y}^{\prime})=\dim\widetilde{Y}^{\prime}.$
@@
$\dim\widetilde{Y}^{\mathrm{Zar}}+\dim u(\widetilde{Y})^{\mathrm{Zar}}<\dim\widetilde{Y}^{\mathrm{ws}}+\dim\widetilde{Y}.$
This contradicts the weak mixed Ax–Schanuel theorem (Theorem 3.5) for $\widetilde{Y}$.
Hence we must have $\widetilde{Y}=\widetilde{Y}^{\prime}$. In particular, $u(\widetilde{Y})$ is algebraic. So $\widetilde{Y}^{\mathrm{Zar}}=\widetilde{Y}^{\mathrm{ws}}$. ∎
For $Y\subseteq S$ irreducible, let $\widetilde{Y}$ be a complex analytic irreducible component of $u^{-1}(Y)$.
@@
If $Y\subseteq X$ is weakly optimal, then $\widetilde{Y}$ is Zariski optimal in $u^{-1}(S)$.
@@
Let $\widetilde{Y}^{\prime}\supseteq\widetilde{Y}$ be a complex analytic irreducible subset of $u^{-1}(S)$ such that $\delta_{\mathrm{Zar}}(\widetilde{Y})\geq\delta_{\mathrm{Zar}}(\widetilde{Y}^{\prime})$. We may and do assume that $\widetilde{Y}^{\prime}$ is Zariski optimal.
Set $Y^{\prime\prime}:=u(\widetilde{Y}^{\prime})^{\mathrm{Zar}}$. Then $Y^{\prime\prime}\subseteq u(\widetilde{Y}^{\prime,\mathrm{ws}})\cap S$ because $u(\widetilde{Y}^{\prime,\mathrm{ws}})\cap S$ is closed and algebraic. Thus
$\delta_{\mathrm{ws}}(Y^{\prime\prime})=\dim Y^{\prime\prime,\mathrm{ws}}-\dim Y^{\prime\prime}\leq\dim\widetilde{Y}^{\prime,\mathrm{ws}}-\dim\widetilde{Y}^{\prime}.$
Since $\widetilde{Y}^{\prime}$ is Zariski optimal, we have $\widetilde{Y}^{\prime,\mathrm{Zar}}=\widetilde{Y}^{\prime,\mathrm{ws}}$ by Theorem 4.2. So, the inequality above further implies
$\delta_{\mathrm{ws}}(Y^{\prime\prime})\leq\dim\widetilde{Y}^{\prime,\mathrm{Zar}}-\dim\widetilde{Y}^{\prime}=\delta_{\mathrm{Zar}}(\widetilde{Y}^{\prime})\leq\delta_{\mathrm{Zar}}(\widetilde{Y})\leq\dim\widetilde{Y}^{\mathrm{ws}}-\dim\widetilde{Y}=\delta_{\mathrm{ws}}(Y).$
So $Y^{\prime\prime}=Y$ because $Y$ is weakly optimal. Thus $\widetilde{Y}=\widetilde{Y}^{\prime}$ is Zariski optimal. ∎
@@ 4.2. A parametrization of Zariski optimal subsets
Fix a Levi decomposition $\mathbf{G}=V\rtimes\mathbf{G}_{0}$. Then for the identification ${\mathscr{D}}=V({\mathbb{R}})\times{\mathscr{D}}_{0}$ in Proposition B.3.(i), the action of $\mathbf{G}({\mathbb{R}})$ on ${\mathscr{D}}$ is given by
$(v,g_{0})\cdot(x_{V},x_{0})=(v+g_{0}x_{V},g_{0}x_{0}).$
@@
    $\{\widetilde{Y}^{\mathrm{Zar}}=\widetilde{Y}^{\mathrm{ws}}:\widetilde{Y}\subseteq u^{-1}(S)\text{ Zariski optimal, with }\widetilde{Y}\cap\mathfrak{F}\not=\emptyset\}$.
@@
Take $(x,H)\in{\mathscr{N}}^{0}$. We wish to prove that $H({\mathbb{R}})x$ equals $\widetilde{Y}^{\mathrm{Zar}}$ for some Zariski optimal $\widetilde{Y}\subseteq u^{-1}(S)$ which passes through $x\in\mathfrak{F}$.
Let $\widetilde{Y}^{\prime}$ be a complex analytic irreducible component of $u^{-1}(S)\cap H({\mathbb{R}})x$ which passes through $x$ with $\dim\widetilde{Y}^{\prime}=d(x,H)$.
Take $\widetilde{Y}\supseteq\widetilde{Y}^{\prime}$ with $\widetilde{Y}\subseteq u^{-1}(S)$ complex analytic irreducible and $\delta_{\mathrm{Zar}}(\widetilde{Y})\leq\delta_{\mathrm{Zar}}(\widetilde{Y}^{\prime})$. We may and do assume that $\widetilde{Y}$ is Zariski optimal. Then $\widetilde{Y}^{\mathrm{Zar}}=\widetilde{Y}^{\mathrm{ws}}$ is a weak Mumford–Tate domain by Theorem 4.2, and hence $\widetilde{Y}^{\mathrm{Zar}}=H^{\prime}({\mathbb{R}})x$ for some $(x,H^{\prime})\in{\mathscr{N}}$. Thus $\delta_{\mathrm{Zar}}(\widetilde{Y})=\delta(x,H^{\prime})$.
Thus $\widetilde{Y}^{\prime}$ is contained in some irreducible component of $H({\mathbb{R}})x\cap H^{\prime}({\mathbb{R}})x$, which by Lemma 4.4.(ii) is $H^{\prime\prime}({\mathbb{R}})x$ for some $(x,H^{\prime\prime})\in{\mathscr{N}}$. Hence $d(x,H)=d(x,H^{\prime\prime})$, so by definition of ${\mathscr{N}}^{0}$ we have $H({\mathbb{R}})x=H^{\prime\prime}({\mathbb{R}})x$. So $H({\mathbb{R}})x\subseteq H^{\prime}({\mathbb{R}})x$. But $\delta(x,H)=\delta(x,H^{\prime})$. So by definition of ${\mathscr{N}}^{0}$ we have $H({\mathbb{R}})x=H^{\prime}({\mathbb{R}})x$, and hence $H({\mathbb{R}})x=\widetilde{Y}^{\mathrm{Zar}}$. Now we have proved that the first set is contained in the second set and the “Moreover” part of the proposition.
Conversely let $\widetilde{Y}\subseteq u^{-1}(S)$ be Zariski optimal with $\widetilde{Y}\cap\mathfrak{F}\not=\emptyset$. By Theorem 4.2, $\widetilde{Y}^{\mathrm{Zar}}=\widetilde{Y}^{\mathrm{ws}}$ is a weak Mumford–Tate domain, and hence equals $H({\mathbb{R}})x$ for some $(x,H)\in{\mathscr{N}}$. We may furthermore choose this $x$ to be a smooth point of $H({\mathbb{R}})x\cap u^{-1}(S)\cap\mathfrak{F}$; then $\dim\widetilde{Y}=d(x,H)$.
@@
Assume there exists $(x,H^{\prime})\in{\mathscr{N}}$ such that $H^{\prime}({\mathbb{R}})x\subsetneq H({\mathbb{R}})x$ and $d(x,H)=d(x,H^{\prime})$. Then $\dim\widetilde{Y}=d(x,H^{\prime})$ and hence $\widetilde{Y}\subseteq H^{\prime}({\mathbb{R}})x$ by analytic continuation. But then $\widetilde{Y}^{\mathrm{Zar}}\subseteq H^{\prime}({\mathbb{R}})x$ by Lemma 4.4.(i). So $\widetilde{Y}^{\mathrm{Zar}}\subseteq H^{\prime}({\mathbb{R}})x\subsetneq H({\mathbb{R}})x=\widetilde{Y}^{\mathrm{Zar}}$, which is impossible. So for any $(x,H^{\prime})\in{\mathscr{N}}$, we have $H^{\prime}({\mathbb{R}})x\subsetneq H({\mathbb{R}})x\Rightarrow d(x,H)>d(x,H^{\prime})$.
Next, assume there exists $(x,H^{\prime})\in{\mathscr{N}}$ such that $H({\mathbb{R}})x\subsetneq H^{\prime}({\mathbb{R}})x$ and $\delta(x,H)\geq\delta(x,H^{\prime})$. Take a complex analytic irreducible component $\widetilde{Y}^{\prime}$ of $u^{-1}(S)\bigcap H^{\prime}({\mathbb{R}})x$ such that $\dim\widetilde{Y}^{\prime}=d(x,H^{\prime})$. Then $\widetilde{Y}\subseteq\widetilde{Y}^{\prime}$ by analytic continuation. By Lemma 4.4.(i), we have $\widetilde{Y}^{\prime,\mathrm{Zar}}\subseteq H^{\prime}({\mathbb{R}})x$. So $\delta_{\mathrm{Zar}}(\widetilde{Y}^{\prime})\leq\delta(x,H^{\prime})\leq\delta(x,H)=\delta_{\mathrm{Zar}}(\widetilde{Y})$. Hence $\widetilde{Y}^{\prime}=\widetilde{Y}$ because $\widetilde{Y}$ is Zariski optimal. So $\delta(x,H^{\prime})=\delta(x,H)$ and $d(x,H^{\prime})=d(x,H)$, and so